# Toward higher-order infinite time Turing machines: simulational $\Gamma$-machines

### v1

@@ 3 General definitions and conditions on operators
In order to formaly study limit rules, we need first to define what a limit rule is. In [Wel00], Welch introduces the concept of operator. A $n$-symbol limit rule can be seen as an operator $\Gamma:{}^{<On}(^{\omega}n)\rightarrow{}^{\omega}n$, that is a function that given any computation history seen as an ordinal-indexed limit sequence of $n$-symbol real numbers (corresponding to tape contents) yields a $n$-symbol real number that will represent the tape content at the next limit stage. As noted, the data of an operator $\Gamma$ is enough to produce a model of infinite time Turing machine whose limit stage behavior is ruled by this operator. Such machines will be called $n$-symbol $\Gamma$-machines. We start by introducing some general concepts.
@@ 3.2 General definitions on machines
Definition 3.8 (Operators and $\Gamma$-machines, [Wel00]).
@@
Definition 3.11 (Writable real numbers, [HL00]).
@@
Definition 3.13 (Eventually writable real numbers, [HL00]).
@@
Definition 3.14 (Accidentally writable real numbers, [HL00]).
@@
Definition 3.16 (Writable ordinals, [HL00]).
@@
Definition 3.17 (Clockable ordinals, [HL00]).
@@
With those definitions, what Hamkins and Lewis asked in [HL00] is whether $\lambda_{\Gamma_{\sup}}=\gamma_{\Gamma_{\sup}}$. As noted, Welch answered this question positively in [Wel00a]. More generally, we may wonder under which conditions on $\Gamma$ does the equality $\lambda_{\Gamma}=\gamma_{\Gamma}$ holds. As done in the discussion of Section 2, we can get a better insight on the question by introducing the following extensions of the concept of clockable, akin to that of the concept of writable.
@@ 3.3.1 Simulational operators
Definition 3.21 (Suitable operator, [Wel00]).
@@
Let us now look at the hypermachines developed by Welch and Friedman in [FW11], under the scrutiny of those new notions. Briefly, the structure of the hypermachine is akin to that of a $2$-symbol ITTM with the difference that it has a new tape called a rule tape. At a limit stage, the values of the cells are computed by a discretionary $\liminf$ operator, that is a $\liminf$ operator that considers only a precise subset of the previous stages over which it takes the $\liminf$. And it is the rule tape which dictates, at a given limit stage, which subset of previous stages should be considered in the $\liminf$.
Now, as this operator needs to look at the whole rule tape, it is not cell-by-cell. However, it is easily seen to be stable. Then, if it was asymptotic, at some limit stage $\alpha\cdot 2$, the limit rule would not depend on the first half of the history. However, the notions of $1$-stability and $1$-correctness, used to define the operator of the hypermachine, are themselves not asymptotic. With those, it is possible to devise an history of computation for the hypermachine which would contradict asymptoticity. Similarly, those definitions are not contraction-proof. Hence, the behavior of the limit rule may change after a contraction or a dilation of the history. A more detail analysis of the hypermachine under those four definitions can be found in [Gir24, Remark 5.3.34].
@@
This is proved as with classical ITTMs. Let $m_{x}$ be a machine which eventually writes $x$ and $m_{y}$ a machine which $x$-writes $y$. With those, we can design a computation which, in parallel, using a classical dovetailing technique, simulates $m_{x}$ and tries to writes $y$ using $m_{y}$ and, as input for $m_{y}$, whatever is written on the output of $m_{x}$. It restarts the simulation of $m_{y}$ each time the output of $m_{x}$ changes. At some point, $m_{x}$ eventually wrote $x$ on its output and this output won’t be modified again. From this point onward, the computation of $m_{y}$ will be conducted using $x$, which will eventually yield $y$. See [Car19, Theorem 3.5.10] and [Win07, Proposition 6.3] for more details. ∎
@@
To show that $\Gamma_{\sup}$ is a suitable $\Sigma_{2}$ operator, it is enough to provide a $\Sigma_{2}$ formula defining the $\limsup$ operator in the sense of Definition 3.21. The only technical part is the fact that this formula does not directly take as argument the history of the machine and that the history is itself defined by this same formula. This can be done in a rather straightforward fashion akin to $\Sigma$-recursion. A more detailed construction of this $\Sigma_{2}$ formula can be found in the proof of Proposition 5.3.37 in [Gir24].
@@ 3.3.2 Looping stability
The possibility for an infinite machine to exit a repeating pattern, as presented in Example 3.26, is in a way what enables those machines to compute for so long and to give rise to intricate computations. At the same time it may lead to pathological patterns from the point of view of those being computation models. As presented in Remark 3.40, this could result in a $\Gamma$-machine which repeats for a very long time (very long when compared to the time it took to enter this repeating pattern) and then suddenly exits this pattern. To avoid those cases, we introduce in Definition 3.39 the notion of looping stability. It is a simple assumption that let us develop a rather general and sound subclass of the simulational operators for which we can prove strong results like Theorem 4.14. Observe that it thus excludes some more exotic models of ITTM, for example cardinal-recognizing ITTMs as introduced in [Hab13].
@@
[…]
@@ 4 Toward higher-order and many-symbol ITTMs
While, as illustrated above, many results on ITTMs are readily generalized to $\Gamma$-machines (provided basic hypothesis on $\Gamma$, e.g. being simulational), there is one interesting result which does not easily lend itself to generalization. As kindly pointed out by Bruno Durand, this is the case of the Speed-up Lemma of [HL00]. This lemma reads: “if $\alpha+k$ is clockable for $k\in\omega$, then so is $\alpha$” and its proof is a clever argument which relies on combinatorial aspects of the $\limsup$ rule. It is clear that this argument can’t be ported as is to $\Gamma$-machines and unclear whether it can be proved differently for those.
Thankfully, the consequences of this fact are limited. One reason is that when it comes to the behavior of infinite machines from a macro perspective (as mostly done here, where the ordinals we consider are limit ordinals closed under many different kind of operations), this lemma is not needed. It is however useful when conducting a finer analysis of gaps and particularly of their beginning and ends. While I reckon most of the analysis regarding the beginning of gaps (see e.g. [BDL23] and [Wel09]) and their links to admissible ordinals may be extended to $\Gamma$-machines, the absence of the Speed-up Lemma makes it less clear when it comes to the end of gaps. In particular, this leads to the following question: does there exist a simulational operator $\Gamma$ for which a gap ends at a non-limit ordinal stage?
@@
We may be tempted, under the hypothesis that $\Gamma$ emulates the operator $\Gamma_{\sup}$, to try and immediatly establish the converse implication of Proposition 4.12 (which will be a corollary of Theorem 4.14). That is to show, for the first case, that if $x$ is accidentally writable then $x\in L_{\Sigma_{\Gamma}}$. It would simply build on the work which was already done for the operator $\Gamma_{\sup}$. Suppose that $x$ is accidentally writable by $\Gamma$, then as $\Gamma$ emulates $\Gamma_{\sup}$ a $\Gamma$-machine could use this $x$ to accidentally write greater and greater ordinals in $\Sigma^{x}_{\Gamma_{\sup}}$ (which is $\Sigma_{\Gamma_{\sup}}$ relativised to computations with $x$ as input). And as we might think that $x\in L_{\Sigma^{x}_{\Gamma_{\sup}}}$ (after all, $x$ is $x$-writable), this would mean, a fortiori, that $x\in L_{\Sigma_{\Gamma}}$. But it happens that $x\in L_{\Sigma^{x}_{\Gamma_{\sup}}}$ simply does not hold in the general case (the fact that $x$ is $x$-writable simply says that $x\in L_{\Sigma^{x}_{\Gamma_{\sup}}}[x]$). Welch first considered in [Wel00] the (analogous) set
@@
Both of those sets are thin (i.e. they do not admit non-empty and non-unit closed subsets) which implies that uncountably many real numbers $x$’s are not in $F_{0}$ or $Q$. A proof of which we give the idea can be found in [Kec75]. Consider $F_{0}$ and suppose that it is not thin. That is, it contains a perfect subset $P$ (i.e. closed subset with not isolated points). $P$ is continuously isomorphic to $^{\omega}2$, so for our purpose we can assume that $P={}^{\omega}2$. We then consider the application $f:x\mapsto\lambda^{x}$. It induces a pre-well-ordering $\preceq$ on $^{\omega}2$:
@@
This theorem was first established in [Wel00a] for the ITTM and, using Theorem 4.14, the proof is the same in the general case. We give one direction of the $\Sigma_{2}$ end-extension.
@@ 5 A counter-example without the looping condition
As for it being $\Delta_{4}$-suitable, the definition provides the skeleton that we can use to define a formula $\varphi_{esc}$ which suits Definition 3.21. The details rest on a recursive definition, as in the proof of Proposition 3.38. A detailed proof with the definition of $\varphi_{esc}$ can be found in [Gir24, Proposition 5.5.9].
@@
As for it being $\Delta_{4}$-suitable, the definition again provides the skeleton to define a formula and Proposition 5.15 ensures that with can distinguish with a formula of bounded complexity whether the limit stage is a multiple of $\nu$. For a complete proof, the reader is referred to [Gir24, Proposition 5.5.16].
@@ 6 Conclusion
Another significant result is the $\lambda$-$\zeta$-$\Sigma$ theorem of [Wel00a]. As done in Corollary 4.33 It can be established in the setting of simulational and looping stable operators using Theorem 4.14. However, this theorem deals with operators definable in set theory with formulas of any complexity. That is, while the $\limsup$ operator is defined by a $\Sigma_{2}$ formula, other operators may be defined by higher-order formulas. And a question is whether, for those operators, we can establish a higher-order equivalent of the $\lambda$-$\zeta$-$\Sigma$ theorem. This was partially answered in the Chapter 6 of the thesis of Johan Girardot. In this chapter, a higher-order operator is designed. It satisfies the constraints which we established here and, for the model of infinite machine it yields, we can generalize the $\lambda$-$\zeta$-$\Sigma$ theorem up to a $\Sigma_{3}$ elementary end-extension.

### v2

@@ 3 General definitions and conditions on operators
In order to formaly study limit rules, we need first to define what a limit rule is. In [16], Welch introduces the concept of operator. A $n$-symbol limit rule can be seen as an operator $\Gamma:{}^{<On}(^{\omega}n)\rightarrow{}^{\omega}n$, that is a function that given any computation history seen as an ordinal-indexed limit sequence of $n$-symbol real numbers (corresponding to tape contents) yields a $n$-symbol real number that will represent the tape content at the next limit stage. As noted, the data of an operator $\Gamma$ is enough to produce a model of infinite time Turing machine whose limit stage behavior is ruled by this operator. Such machines will be called $n$-symbol $\Gamma$-machines. We start by introducing some general concepts.
@@ 3.2 General definitions on machines
Definition 3.8 (Operators and $\Gamma$-machines, [16]).
@@
Definition 3.11 (Writable real numbers, [12]).
@@
Definition 3.13 (Eventually writable real numbers, [12]).
@@
Definition 3.14 (Accidentally writable real numbers, [12]).
@@
Definition 3.16 (Writable ordinals, [12]).
@@
Definition 3.17 (Clockable ordinals, [12]).
@@
With those definitions, what Hamkins and Lewis asked in [12] is whether $\lambda_{\Gamma_{\sup}}=\gamma_{\Gamma_{\sup}}$. As noted, Welch answered this question positively in [17]. More generally, we may wonder under which conditions on $\Gamma$ does the equality $\lambda_{\Gamma}=\gamma_{\Gamma}$ holds. As done in the discussion of Section 2, we can get a better insight on the question by introducing the following extensions of the concept of clockable, akin to that of the concept of writable.
@@ 3.3.1 Simulational operators
Definition 3.21 (Suitable operator, [16]).
@@
Let us now look at the hypermachines developed by Welch and Friedman in [9], under the scrutiny of those new notions. Briefly, the structure of the hypermachine is akin to that of a $2$-symbol ITTM with the difference that it has a new tape called a rule tape. At a limit stage, the values of the cells are computed by a discretionary $\liminf$ operator, that is a $\liminf$ operator that considers only a precise subset of the previous stages over which it takes the $\liminf$. And it is the rule tape which dictates, at a given limit stage, which subset of previous stages should be considered in the $\liminf$.
Now, as this operator needs to look at the whole rule tape, it is not cell-by-cell. However, it is easily seen to be stable. Then, if it was asymptotic, at some limit stage $\alpha\cdot 2$, the limit rule would not depend on the first half of the history. However, the notions of $1$-stability and $1$-correctness, used to define the operator of the hypermachine, are themselves not asymptotic. With those, it is possible to devise an history of computation for the hypermachine which would contradict asymptoticity. Similarly, those definitions are not contraction-proof. Hence, the behavior of the limit rule may change after a contraction or a dilation of the history. A more detail analysis of the hypermachine under those four definitions can be found in [10, Remark 5.3.34].
@@
This is proved as with classical ITTMs. Let $m_{x}$ be a machine which eventually writes $x$ and $m_{y}$ a machine which $x$-writes $y$. With those, we can design a computation which, in parallel, using a classical dovetailing technique, simulates $m_{x}$ and tries to writes $y$ using $m_{y}$ and, as input for $m_{y}$, whatever is written on the output of $m_{x}$. It restarts the simulation of $m_{y}$ each time the output of $m_{x}$ changes. At some point, $m_{x}$ eventually wrote $x$ on its output and this output won’t be modified again. From this point onward, the computation of $m_{y}$ will be conducted using $x$, which will eventually yield $y$. See [1, Theorem 3.5.10] and [18, Proposition 6.3] for more details. ∎
@@
To show that $\Gamma_{\sup}$ is a suitable $\Sigma_{2}$ operator, it is enough to provide a $\Sigma_{2}$ formula defining the $\limsup$ operator in the sense of Definition 3.21. The only technical part is the fact that this formula does not directly take as argument the history of the machine and that the history is itself defined by this same formula. This can be done in a rather straightforward fashion akin to $\Sigma$-recursion. A more detailed construction of this $\Sigma_{2}$ formula can be found in the proof of Proposition 5.3.37 in [10].
@@ 3.3.2 Looping stability
The possibility for an infinite machine to exit a repeating pattern, as presented in Example 3.26, is in a way what enables those machines to compute for so long and to give rise to intricate computations. At the same time it may lead to pathological patterns from the point of view of those being computation models. As presented in Remark 3.40, this could result in a $\Gamma$-machine which repeats for a very long time (very long when compared to the time it took to enter this repeating pattern) and then suddenly exits this pattern. To avoid those cases, we introduce in Definition 3.39 the notion of looping stability. It is a simple assumption that let us develop a rather general and sound subclass of the simulational operators for which we can prove strong results like Theorem 4.14. Observe that it thus excludes some more exotic models of ITTM, for example cardinal-recognizing ITTMs as introduced in [11].
@@
Consider ordinal $\omega_{2}$. By the downward Löwenheim-skolem theorem, we can find limit ordinals $\alpha_{1}<\alpha_{2}<\alpha_{3}$ of cardinality $\aleph_{1}$ and such that the $L_{\alpha_{i}}$ are elementary substructures of $L_{\omega_{2}}$. Moreover we can arrange for $\alpha_{3}$ to be strictly greater than $\alpha_{2}\cdot\omega$. Further, by a result of Devlin [7, II. 5.5], we can show in $L$ that no new real numbers appear in the constructible universe after construction stage $\omega_{1}$. Hence, we have:
@@ 4 Toward higher-order and many-symbol ITTMs
While, as illustrated above, many results on ITTMs are readily generalized to $\Gamma$-machines (provided basic hypothesis on $\Gamma$, e.g. being simulational), there is one interesting result which does not easily lend itself to generalization. As kindly pointed out by Bruno Durand, this is the case of the Speed-up Lemma of [12]. This lemma reads: “if $\alpha+k$ is clockable for $k\in\omega$, then so is $\alpha$” and its proof is a clever argument which relies on combinatorial aspects of the $\limsup$ rule. It is clear that this argument can’t be ported as is to $\Gamma$-machines and unclear whether it can be proved differently for those.
Thankfully, the consequences of this fact are limited. One reason is that when it comes to the behavior of infinite machines from a macro perspective (as mostly done here, where the ordinals we consider are limit ordinals closed under many different kind of operations), this lemma is not needed. It is however useful when conducting a finer analysis of gaps and particularly of their beginning and ends. While I reckon most of the analysis regarding the beginning of gaps (see e.g. [5] and [3]) and their links to admissible ordinals may be extended to $\Gamma$-machines, the absence of the Speed-up Lemma makes it less clear when it comes to the end of gaps. In particular, this leads to the following question: does there exist a simulational operator $\Gamma$ for which a gap ends at a non-limit ordinal stage?
@@
We may be tempted, under the hypothesis that $\Gamma$ emulates the operator $\Gamma_{\sup}$, to try and immediatly establish the converse implication of Proposition 4.12 (which will be a corollary of Theorem 4.14). That is to show, for the first case, that if $x$ is accidentally writable then $x\in L_{\Sigma_{\Gamma}}$. It would simply build on the work which was already done for the operator $\Gamma_{\sup}$. Suppose that $x$ is accidentally writable by $\Gamma$, then as $\Gamma$ emulates $\Gamma_{\sup}$ a $\Gamma$-machine could use this $x$ to accidentally write greater and greater ordinals in $\Sigma^{x}_{\Gamma_{\sup}}$ (which is $\Sigma_{\Gamma_{\sup}}$ relativised to computations with $x$ as input). And as we might think that $x\in L_{\Sigma^{x}_{\Gamma_{\sup}}}$ (after all, $x$ is $x$-writable), this would mean, a fortiori, that $x\in L_{\Sigma_{\Gamma}}$. But it happens that $x\in L_{\Sigma^{x}_{\Gamma_{\sup}}}$ simply does not hold in the general case (the fact that $x$ is $x$-writable simply says that $x\in L_{\Sigma^{x}_{\Gamma_{\sup}}}[x]$). Welch first considered in [16] the (analogous) set
@@
Both of those sets are thin (i.e. they do not admit non-empty and non-unit closed subsets) which implies that uncountably many real numbers $x$’s are not in $F_{0}$ or $Q$. A proof of which we give the idea can be found in [2]. Consider $F_{0}$ and suppose that it is not thin. That is, it contains a perfect subset $P$ (i.e. closed subset with not isolated points). $P$ is continuously isomorphic to $^{\omega}2$, so for our purpose we can assume that $P={}^{\omega}2$. We then consider the application $f:x\mapsto\lambda^{x}$. It induces a pre-well-ordering $\preceq$ on $^{\omega}2$:
@@
This theorem was first established in [17] for the ITTM and, using Theorem 4.14, the proof is the same in the general case. We give one direction of the $\Sigma_{2}$ end-extension.
@@ 5 A counter-example without the looping condition
As for it being $\Delta_{4}$-suitable, the definition provides the skeleton that we can use to define a formula $\varphi_{esc}$ which suits Definition 3.21. The details rest on a recursive definition, as in the proof of Proposition 3.38. A detailed proof with the definition of $\varphi_{esc}$ can be found in [10, Proposition 5.5.9].
@@
As for it being $\Delta_{4}$-suitable, the definition again provides the skeleton to define a formula and Proposition 5.15 ensures that with can distinguish with a formula of bounded complexity whether the limit stage is a multiple of $\nu$. For a complete proof, the reader is referred to [10, Proposition 5.5.16].
@@ 6 Conclusion
Another significant result is the $\lambda$-$\zeta$-$\Sigma$ theorem of [17]. As done in Corollary 4.33 It can be established in the setting of simulational and looping stable operators using Theorem 4.14. However, this theorem deals with operators definable in set theory with formulas of any complexity. That is, while the $\limsup$ operator is defined by a $\Sigma_{2}$ formula, other operators may be defined by higher-order formulas. And a question is whether, for those operators, we can establish a higher-order equivalent of the $\lambda$-$\zeta$-$\Sigma$ theorem. This was partially answered in the Chapter 6 of the thesis of Johan Girardot. In this chapter, a higher-order operator is designed. It satisfies the constraints which we established here and, for the model of infinite machine it yields, we can generalize the $\lambda$-$\zeta$-$\Sigma$ theorem up to a $\Sigma_{3}$ elementary end-extension.